OpenStudy (anonymous):

Please solve 1.)4 − 3x = 7x − 6 2.)8 + 4x = −10 + x

OpenStudy (anonymous):

Move the constants to one side and the varibles to another, then divide by the number in front of the varible. \[ 4-3x=7x-6\] \[ -3x - 7x = -6 - 4\] You should be able to take it from there

OpenStudy (anonymous):

-3x-7x=-6-4......transpose 7x to other side and 4 -10x=-10.....add the like terms x=1.....divide both side by -10 and get your answer

OpenStudy (anonymous):

4 − 3x = 7x − 6 Put x on right and the rest on the other part(i mean afte= ) 4+6=7x+3x 10=10x x=10/10 x=1

OpenStudy (anonymous):

after*

OpenStudy (anonymous):

8 + 4x = −10 + x same will be here 4x-x=-10-8 3x=-18 x=-18/3 x=-6

OpenStudy (anonymous):

Thank you!! I wrote that all down so I could actually understand the problem

OpenStudy (anonymous):

Hahaha okay if you don't understand I'll explain you once again :)

OpenStudy (anonymous):

We just have to find x :)

OpenStudy (anonymous):

8+4x=-10+x 4x-x=-10-8 3x=-18.....divide both sides by 3 x=-6..